OpenStudy (anonymous):

Using Jack's simulation, what is the probability of rolling a 6 on the number cube and the coin landing on heads? fraction 1,107 over 10,000 fraction 1,593 over 10,000 fraction 27 over 100 fraction 41 over 100

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

@Lexi_Loves @linksonic1 @LanHikari22

OpenStudy (smd_kma):

fraction 1,107 over 10,000

OpenStudy (anonymous):

How do you know that?

OpenStudy (anonymous):

How did you figure it out?

OpenStudy (anonymous):

hello?

OpenStudy (benlindquist):

yes

OpenStudy (anonymous):

How do you know it is fraction 1,107 over 10,000?

OpenStudy (anonymous):

@nuccioreggie If you help me ill help you

OpenStudy (anonymous):

@pooja195 @misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

would help if we saw the simulation right?

OpenStudy (anonymous):

Hello!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Number on the Cube Number of Times Rolled 1 16 2 14 3 5 4 17 5 21 6 27 Heads Tails 41 59

OpenStudy (anonymous):

@misty1212 please come back

OpenStudy (misty1212):

the total is 49+51=100 right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

lolol this must be some other simulation it has heads and tails, like a coin, but your question is about rolling a 6 on a die i think you must be real confused by this

OpenStudy (misty1212):

oooh i see nvm i was confused we can do this

OpenStudy (anonymous):

this is the dice part: Number on the Cube Number of Times Rolled 1 16 2 14 3 5 4 17 5 21 6 27

OpenStudy (anonymous):

ok lol

OpenStudy (misty1212):

head is \(\frac{49}{100}\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OH WAIT IT IS 41

OpenStudy (misty1212):

how about 6's?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it is 27/100

OpenStudy (misty1212):

ok heads is \(\frac{41}{100}\) how about sixes/?

OpenStudy (misty1212):

ok then your last job is to multiply those two fractions

OpenStudy (anonymous):

\[\frac{27}{100}\]

OpenStudy (anonymous):

ok tyhanks\

OpenStudy (misty1212):

\[\frac{41}{100}\times \frac{27}{100}\]

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

\[\frac{41}{100}\times \frac{27}{100}\]

OpenStudy (anonymous):

is 1107 as the numerator and 10000 as the denominator

OpenStudy (anonymous):

yay thanks!